An industry provides 600000 m3 fermenter for the production of single 1/1 cell protein. When the media for SCP is prepared, it anticipated that it contains about 10^(7) organisms per cm3 of the fermenter. Determine the del factor for the system.
1
Expert's answer
2021-09-20T14:52:02-0400

∇ = ln ("\frac{N_{0}}{N_{t}}" )


"∇ = ln (\frac{600000}{10000000} )"


"\u2207 = ln (0.06)"


"\u2207 = -2.8134"